Consider the equations 5y2=16x and y2=8x-24. What is the pointbof intersection of the given curves on the first quadrant?=

1
Expert's answer
2021-05-21T01:14:03-0400

"\text{Given, two curves}\space
 5y^2=16x \space and \space y^2=8x-24.\newline
Let f_1:y=\sqrt{\frac{16x}{5}} and f_2:y=\sqrt{8x-5}.\newline
\text{Point of interestion are those points where the given two graph equation have same value.}\newline
\text{Intersecting points given by,}
f_1=f_2\newline
\sqrt{\frac{16x}{5}}=\sqrt{8x-5}\newline
\frac{16x}{5}=8x-24\newline
16x=40x-120\newline
x=5\newline
Then,
 y=\sqrt{8×5-24}\newline
=\pm4
\text{Therefore, point of interestion are }(5,-4)and(5,4)."